Question 42 of 45 (1 point) For what value of r is the statement an identity? -X-9 = x + 5+ provided that x #6 X-6 X-6 r =
Answer #1

32-x-9 = x+5 + r De-6 X-6 solution- sc²- x-9 – (x+5) C-6 Se-6 10 x²-x-9 - (x+5) (x-6) x?-x-9 - {x?-6x+58-30? = or { x²-x-31 -